OpenStudy (anonymous):

If a function is defined as f(x)=x^2 then translated downward 4 units and over to the right 3 units what is h(x)?

jimthompson5910 (jim_thompson5910):

Translate downward 4 units: f(x) - 4 = x^2 - 4 Translate to the right 3 units: f(x-3) - 4 = (x-3)^2 - 4 So h(x) = (x-3)^2 - 4

OpenStudy (anonymous):

wait no, i had it backwards. over to the left 3 units and up four

jimthompson5910 (jim_thompson5910):

Then it's h(x) = (x+3)^2 + 4

OpenStudy (anonymous):

ohhh, i getcha:)